### v2

@@ 4 Experimental Design
We conduct an extensive series of over 150 experiments on the more than 35 models described in Section 3.3 to evaluate which models are ideal under a given set of constraints. We measure efficiency w.r.t. the full models rather than individual attention modules, since methods that reduce the token sequence (Section 3.3 (ii)), and others, like XCiT, require additional modifications to the architecture, which we include. Note, however, that many models only introduce changes in the attention mechanism, effectively turning the benchmark into a module-level comparison for these.
@@ 4.1 Baselines
To find the most efficient models, we need to quantify their efficiency gains by comparing them to the original ViT [14], and to its follow-up version [50, 51]. We also include a ResNet50 [19] in our evaluation as a representative baseline for CNN architectures, and a point of reference across papers. All evaluations are based on the ImageNet-1k dataset for classification [13], as it is one of the most commonly used for benchmarks in this domain. Further results for metrics that run fine-tuning on other datasets can be found in the supplementary material.
@@ 4.2 Training Pipeline
We compare models on even grounds by training using a standardized pipeline. To reduce bias, our pipeline is relatively simple and only consists of elements commonly used in CV. In particular, we refrain from using knowledge distillation to prevent introducing biases from the choice of teacher model. Any orthogonal techniques, like quantization, sample selection, and others, are not included as they can be applied to every model and would manifest as a systematic offset in the results. To avoid issues stemming from limited training data, we pre-train all models on ImageNet-21k [38].
Models are trained for a total of 140 epochs using the training pipeline introduced in [51]. This is an updated version of the pipeline used by DeiT [50], which has been adopted throughout the literature for training efficient ViTs (see Table 1). As this pipeline only employs standard CV practices, we consider it a fair point of comparison for evaluating all models.
Pretraining is conducted on ImageNet-21k for 90 epochs at resolutions of $224$ and $192$ pixels, followed by fine-tuning on ImageNet-1k at $224$ and $384$ pixels for 50 epochs. In cases where model training is unstable, we default back to the hyperparameters reported in the corresponding publications. All training is conducted using 4 or 8 NVIDIA A100 GPUs. While most of the models work well with our training pipeline, we could not get Performer, Linear Transformer, or HaloNet to converge. See the supplementary material for more details.
@@ 4.3 Efficiency Metrics
As we saw, efficiency has multiple dimensions, and therefore we need to capture a variety of metrics. For comparison with other papers and to evaluate their use as proxy metrics for speed and memory requirements, we capture the theoretical metrics of number of parameters and FLOPS. These are used by the community to provide estimates of the model’s representational capacity and computational requirements [35, 48, 37]. However, theoretical metrics do not always correlate with real-world performance, especially when comparing different architectures [12, 4]. Instead, empirical metrics, obtained by running models on hardware, offer more informative evaluations, which is why we track GPU time for training, inference throughput at optimal batch sizes, and VRAM requirements. While these empirical metrics can be sensitive to hardware and software configurations, we ensure consistent evaluations by employing the same setup for all models [^2] We evaluate the variance of our measurements in the supplementary material.
@@ 5 Results
Table 1: ImageNet-1k accuracy of the original papers and the new training pipeline. ✓ marks original pipelines based on DeiT [50]. Models are trained on $224$px images, unless marked with $\uparrow$ ($384$px) or $\downarrow$ ($112$px). Results with ${}^{\star}$ use knowledge distillation.
@@
Model Original Ours
DeiT Accuracy Accuracy
@@
FocalNet-S $83.4$ $84.91$
DynamicViT-S ${}^{\star}$ $81.09$
EViT (delete) $79.4$ $82.29$
EViT (fuse) $79.5$ $81.96$
@@ 5.1 Improved Training Pipeline
To validate the fairness of our training pipeline, in Table 1 we compare ImageNet-1k accuracy reported in the original papers (whenever reported) with the one we obtained, to see if there are any obvious outliers. Firstly, note that a substantial fraction (9 out of 19) of papers base their training pipelines on DeiT [50], making them a good fit for training with the updated version from DeiT III [51].
We observe that almost all the models trained with this pipeline achieve a higher accuracy; 1.35% on average, with the largest gap being +6.86% for Poly-SA. Note that the only three models reporting higher performance than what we obtain use either knowledge distillation, which we avoid to not induce bias, or fine-tune at a higher resolution, which as we will see later is inefficient.
Most categories contain models that achieve peak accuracy of approximately 85% (see the supplementary material). However, two notable exceptions are observed: the Kernel Attention class, which poses challenges in optimizing models, and the Fixed Attention class, where the use of a constant attention matrix inherently results in a lower accuracy. Overall, our selection of the training pipeline as a reference for a fair comparison is supported by the improvements in results compared to the original papers, regardless of architectural differences.
Figure 3: Accuracy (red line, right y-axis) and accuracy per parameter (bars, left y-axis) of different models at a resolution of $224\times 224$ pixels. Models are ordered by accuracy.
Figure 4: Pareto front (dotted line) of number of parameters vs. accuracy for models with at most 30 million parameters. Models using higher image resolutions dominate, since their drawbacks are not represented for this metric.
@@ 5.2 Number of Parameters
The number of parameters is used in the literature as a proxy metric for tracking the model complexity, and the overall computational cost of using a model. When analyzing the parameter efficiency of different models (Figure 3), it is evident that for most smaller models, the accuracy per parameter remains relatively constant at a baseline level of about $4\times 10^{-8}\frac{\%}{\text{param.}}$. This value is approximately halved for larger models, indicating diminishing returns on scaling the model size. For ViT, there is a noticeable drop in accuracy per parameter across the model sizes. While ViT-Ti outperforms models of comparable accuracy, and ViT-S performs on par with similar models, ViT-B slightly underperforms when compared to other, larger models. Significant outliers are the smaller Hybrid Attention models EfficientFormerV2-S0 and CoaT-Ti, which exhibit the highest accuracy per parameter, outperforming other attention-based models as well as ResNet50. This suggests that the combination of attention and convolutions allows for the development of very parameter-efficient models.
Figure 4 highlights the Pareto boundary between accuracy and the number of parameters, showing models with fewer than $30$ million parameters. It is noteworthy that the majority of Pareto optimal models are fine-tuned at a higher resolution of $384$ pixels, which contributes to an increase in accuracy without any growth in the parameter count. Here, the number of parameters does not accurately represent the gain in computational cost. At the smallest model sizes, we observe once again, that models relying on Hybrid Attention e.g., EfficientFormerV2-S0 and CoaT-Ti are Pareto optimal. The TokenLearner is a Pareto optimal choice between ViT-Ti and the also Pareto optimal ViT-S.
@@ 5.3 Speed
Whether driven by strict requirements for real-time processing or the desire to obtain model outputs within reasonable timeframes, inference speed directly impacts the usability of deployed models. The models we evaluate often claim a superior throughput vs. accuracy trade-off compared to ViT. However, our evaluation (Figure 1) reveals that ViT remains Pareto optimal at all sizes. Only few models, namely Synthesizer-FR and some Sequence Reduction models, show improvements in the Pareto front when compared to a ViT of comparable size.
We also observe that most models maintain their Pareto optimality even when fine-tuned on other datasets like Stanford Cars [24], Oxford Flowers 102 [33], and MIT Places365 [69]. Even when assessing throughput on a CPU instead of a GPU, we observe only minimal fluctuations in model scaling. See supplementary material for more information about these results.
Our observations reveal that fine-tuning at a higher resolution is inefficient. While it may result in improved accuracy, it also entails a significant increase in computational cost, leading to a substantial reduction in throughput. In turn, scaling up the model ends up being more efficient. This is shown in Figure 1, where no model fine-tuned at the higher resolution is Pareto optimal. For example, when scaling from ViT-Ti to ViT-S (points A to C), we gain a lot more accuracy for the lost speed than when scaling up the image size (points A to B).
In our analysis of fine-tuning time, we see many similarities to the Pareto front of inference time. Additionally, the TokenLearner model emerges as a standout performer with the fastest fine-tuning speed while achieving a competitive accuracy of 77.35%. For more details, refer to the supplementary material.
@@ 5.4 Memory
Figure 5 (right) reveals that the tradeoffs when optimizing for VRAM usage during inference are very different from the ones for speed or training VRAM (left side), making it the only metric where ViT is not Pareto optimal. At inference time, the Hybrid Attention models CoaT and CvT stand out from the rest. Additionally, CaiT, and Wave-ViT, and in contrast to the other metrics the Sequence Reduction models EViT and ToMe, fine-tuned at $384$ px, form the Pareto front for larger sizes. This observation suggests that, similar to the number of parameters, Hybrid Attention models excel in low memory environments.
In contrast, the aspect of training memory (Figure 5 (left)) exhibits a similar pattern as observed in throughput. Here, CoaT and CvT need more memory than other models of comparable accuracy, and again ViT is Pareto optimal, along with the Sequence Reduction models EViT, ToMe, TokenLearner, and Synthesizer-FR, and XCiT all at $224$px. When training at a resolution of $384$px most models need more than $225$GB of VRAM and therefore are not plotted.
@@ 5.5 Correlation of Metrics
We run a correlation analysis between the efficiency metrics in Table 2. The highest correlation of $0.89$ is between fine-tuning time and training memory. This suggests a common underlying factor or bottleneck, possibly related to the necessity of memory reads during training. Understanding this relationship can provide valuable insights into the factors influencing training efficiency. Intriguingly, the highest correlation coefficient involving a theoretical metric is $0.85$, found between FLOPS and training memory, suggesting that the VRAM required for training can be roughly estimated based on the theoretical FLOPS using the following approximation (plot in the supplementary material):
[…]
(FiXme) FiXme Warning: ’Reviewer #2: Relationship might be useful, although evidence is relatively weak’
Surprisingly, the other evaluated metrics exhibit relatively weak correlations when considering different model architectures, highlighting the limited reliability of estimating computational costs solely based on theoretical metrics. Consequently, assessing model efficiency in practical scenarios requires the measurement of throughput and memory requirements for novel architectures.
@@ 6 Discussion & Conclusion
ptWhich Transformerto Favor?pt1. MainConstraintpt2. Phasept3. Transformers to FavorSpeedMemoryInferenceTrainingInferenceTraining{ViT, ToMe, TokenLearner, …}{ViT, EViT, TokenLearner, …}{ViT, CoaT, CvT, WaveViT}{ViT, ToMe, TokenLearner, …}
Figure 6: Flowchart for answering the question Which Transformer to Favor? in image classification. First, choose the main constraint (1.), speed, or memory, and the phase you want to optimize this constraint at (2.): Inference or training. Based on that, we provide recommendations on which transformer to favor (3.) from the taxonomy classes of Baseline, Sequence Reduction, Hybrid Attention, and Sparse Attention. See the supplementary material for a full list of Pareto optimal models for each metric.
In this paper, we provide a thorough benchmark of efficiency oriented transformers for vision, facilitating a fair, competitive comparison of contributions from NLP and vision and revealing some unexpected trends. Results show ViT’s continued Pareto optimality across multiple metrics, reaffirming its effectiveness as a baseline. Results prove that scaling the model size is, in most cases, more efficient than using higher resolution images. When comparing different efficiency metrics, we uncover the limitations of estimating computational costs solely based on theoretical metrics, with an exception being training VRAM requirements, which can be roughly estimated based on the theoretical FLOPS.
When addressing the question of which transformer to favor, our benchmark offers actionable insights in the form of models and strategies to use (see Figure 6). ViT remains the preferred choice overall. However, token sequence methods can become viable alternatives when speed and training efficiency are of importance. Additionally, for scenarios with significant inference memory constraints, considering hybrid attention models can prove advantageous.
Looking ahead, we posit that the establishment of such a comparable, fair benchmark and analysis will drive substantial advancements in efficiency-oriented transformers, benefitting both researchers and practitioners making architectural decisions for resource-constrained environments.

### v3

@@ 4 Experimental Design
We conduct an extensive series of over 200 experiments on the more than 45 models described in Section 3.3 to evaluate which models are ideal under a given set of constraints. We measure efficiency w.r.t. the full models rather than individual attention modules, since methods that reduce the token sequence (Section 3.3 (ii)), and others, like XCiT, require additional modifications to the architecture, which we include. Note, however, that many models only introduce changes in the attention mechanism, effectively turning our benchmark into a module-level comparison for these.
@@ 4.1 Baselines
To find the most efficient models, we need to quantify their efficiency gains by comparing them to the original ViT [16] and its follow-up versions [57, 58]. We also include ResNet50 [24] in our evaluation as a representative baseline for CNN architectures and a point of reference across papers. All evaluations are based on the ImageNet-1k dataset for classification [15] as it is one of the most commonly used datasets in this domain. Results on other datasets (Stanford Cars [28], Oxford Flowers 102 [39], MIT Places365 [79]) can be found in the supplementary material.
@@ 4.2 Training Pipeline
We compare models on even grounds by training with a standardized pipeline. To reduce bias, our pipeline is relatively simple and only consists of elements commonly used in CV. In particular, we refrain from using knowledge distillation to prevent introducing bias from the choice of teacher model. Any orthogonal techniques, like quantization, sample selection, and others, are not included as they can be applied to every model and would manifest as a systematic offset in the results. To avoid issues from limited training data, we pre-train all models on ImageNet-21k [44].
Models are trained for a total of 140 epochs using the pipeline introduced in [58]. This is an updated version of DeiT [57], which has been adopted throughout the literature for training efficient ViTs (see Table 1). As this pipeline only employs standard CV practices, we consider it a fair point of comparison for evaluating all models.
Pretraining is conducted on ImageNet-21k for 90 epochs at resolutions of $224$ and $192$ pixels, followed by fine-tuning on ImageNet-1k at $224$ and $384$ pixels for 50 epochs. In cases where model training is unstable, we default back to the hyperparameters reported in the corresponding publications. All training is conducted using 4 or 8 NVIDIA A100 GPUs. While most of the models work well with our training pipeline, some did not converge. See the supplementary material for more details.
@@ 4.3 Efficiency Metrics
As we saw, efficiency has multiple dimensions, and therefore we need to capture a variety of metrics. For comparison with other papers and to evaluate their use as proxy metrics for speed and memory requirements, we capture the theoretical metrics of number of parameters and FLOPS. These are used by the community to provide estimates of the model’s representational capacity and computational requirements [41, 55, 43]. However, theoretical metrics do not always correlate with real-world performance, especially when comparing different architectures [14, 4]. Instead, empirical metrics, obtained by running models on hardware, offer more informative evaluations, which is why we track GPU time for training, inference throughput at optimal batch sizes, and VRAM requirements. While these empirical metrics can be sensitive to hardware and software configurations, we ensure consistent evaluations by employing the same setup for all models. We evaluate the variance of our metrics in the supplementary material.
@@ 5 Results
Table 1: ImageNet-1k accuracy of the original papers and the new training pipeline. ✓ marks original pipelines based on DeiT [57]. Models are trained on $224$px images, unless marked with $\uparrow$ ($384$px) or $\downarrow$ ($112$px). Results with D use knowledge distillation and E uses exponential moving average.
@@
EfficientViT-B2 82.7$\uparrow$ 81.52
NextViT-S 82.5 83.92
ResT-S 79.6 79.92
FocalNet-S 83.4 84.91
SwiftFormer-S 78.5D 76.41
FastViT-S12 79.8$\uparrow$ 78.77
EfficientMod-S 81.0 80.21
@@
EViT 79.4 82.29
DynamicViT-S 83.0D 81.09
EViT Fuse 79.5 81.96
@@ 5.1 Improved Training Pipeline
To validate the fairness of our training pipeline, in Table 1 we compare ImageNet-1k accuracy reported in the original papers (whenever reported) with the one we obtained to see if there are any obvious outliers. Note, that a substantial fraction (13 out of 26) of papers explicitely base their training pipelines on DeiT [57], making them a good fit for training with the updated version in DeiT III [58]. We observe that almost all models trained with this pipeline achieve a higher accuracy; 0.85% on average, with the largest gain being +6.86% for Poly-SA. The models reporting higher performance than what we obtain use knowledge distillation, which we avoid to not induce bias, or a higher resolution, which we find is inefficient.
Most categories contain models that achieve peak accuracy of approximately 85% (see the supplementary material). However, two notable exceptions are observed: the Kernel Attention class, which might pose challenges in optimizing models, and the Fixed Attention class, where the use of a constant attention matrix inherently results in a lower accuracy. Overall, our selection of the training pipeline as a reference for a fair comparison is supported by the improvements in results compared to the original papers, regardless of architectural differences.
Figure 3: Accuracy (red line, right y-axis) and accuracy per parameter (bars, left y-axis) of models ordered by accuracy at a resolution of $224$px. 14 of 47 models of intermediate size are grouped into others. See the supplementary material for the full plot.
@@ 5.2 Number of Parameters
The number of parameters is used in the literature as a proxy metric for tracking the model complexity and the overall computational cost of using a model. When analyzing the parameter efficiency of different models (Figure 3), it is evident that for most smaller models, the accuracy per parameter remains relatively constant at a baseline level of about $4\times 10^{-8}\frac{\%}{\text{param.}}$. This value is approximately halved for larger models, indicating diminishing returns on scaling the model size. For ViT, there is a noticeable drop in accuracy per parameter across the model sizes. While ViT-Ti outperforms models of comparable accuracy, and ViT-S performs on par with similar models, ViT-B slightly underperforms when compared to other, larger models. Significant outliers are the smaller Hybrid Attention models EfficientFormerV2-S0 and CoaT-Ti, which exhibit the highest accuracy per parameter, as well as those Non-attention Shuffling models which incorporte convolutions, like SwiftFormer and FastViT. In the MLP category, we find HiViT to have significantly higher and Switch ViT with significantly lower accuracy per parameter than surrounding models. This suggests that the combination of attention and convolutions allows for the development of highly parameter-efficient models.
@@ 5.3 Speed
Figure 4: Pareto front of finetuning time and accuracy for models which need less than 50 hours for finetuning. We include the full plot in the supplementary material.
Whether driven by strict requirements for real-time processing or the desire to obtain model outputs within reasonable timeframes, inference speed directly impacts the usability of deployed models. The models we evaluate often claim a superior throughput vs. accuracy trade-off compared to ViT. However, our evaluation in Figure 1 reveals that ViT remains Pareto optimal at all model sizes. Only few models, namely Synthesizer-FR, NextViT, and some Sequence Reduction models, show improvements in the Pareto front when compared to a ViT of comparable size.
We also analyze the Pareto optimality with models finetuned on Stanford Cars [28], Oxford Flowers 102 [39], and MIT Places365 [79] in the supplementary material, observing a high Spearman correlation between the accuracy of different datasets. In particular, the Pareto optimal models of throughput and accuracy are largely the same, which is a strong indicator for the generalizability of our results. Even when assessing throughput on a CPU instead of a GPU, we observe only a high Spearman correlation with only few outliers. Again, we find the Pareto front to be robust to the choice of hardware. See the supplementary material for more information about these results.
In the Pareto front of fine-tuning time in Figure 4, we see many similarities to the one of inference time. Here, some of the token sequence based models are highly efficient. In particular, TokenLearner emerges as a standout performer with the fastest fine-tuning speed while achieving a competitive accuracy of 77.35%. For larger models, we find ViT-B and FocalNet (-Ti and -S) to be Pareto optimal.
Generally, ViT is still a solid choice for speed, with some Token Sequence models trading off accuracy for speed.
@@ 5.4 Memory
Figure 5 (right) reveals that the tradeoffs when optimizing for VRAM usage during inference are very different from the ones for speed or training VRAM (left side), making it the only metric where ViT is not Pareto optimal. At inference time, the Hybrid Attention models CoaT, CvT, ResT, and NextViT stand out from the rest. Additionally, CaiT, and EViT@384px for the larger sizes and EfficientMod for the smaller ones are Pareto optimal. This observation suggests that, similar to the number of parameters, Hybrid Attention models excel in low memory environments.
In contrast, the aspect of training memory (Figure 5; left) exhibits a similar pattern as observed in throughput. Here, CoaT, CvT, and ResT need more memory than other models of comparable accuracy and again ViT is Pareto optimal, along with some Sequence Reduction models and Synthesizer-FR. NextViT is the only model that is Pareto optimal for both training and inference memory. When training at a resolution of $384$px most models need more than $225$GB of VRAM and therefore are not plotted.
Overall, inference memory is the most different from the other Pareto fronts with Hybrid Attention models excelling.
5.5 Scaling Behaviors
Our observations reveal that fine-tuning at a higher resolution is inefficient. While it may result in improved accuracy, it entails a significant increase in computational cost, leading to a substantial reduction in throughput. In turn, scaling up the model ends up being more efficient. This is shown in Figure 1, where the best models using high resolution images are two to four times slower than a low resolution model at the same accuracy (compare Pareto front and HR Pareto front). Consequently, no model fine-tuned at the higher resolution is Pareto optimal. For example, when scaling from tiny to small (A to C in Figure 1), we gain a lot more accuracy for the lost speed than when scaling up the image size (A to B). For training memory, this trend is even more pronounced (see HR Pareto front in Figure 5), with most of the high resolution models requiring more than $225$GB of VRAM for training and thus being off the chart. Overall, many papers focus on reducing the $\mathcal{O}(N^{2})$ complexity of self-attention to deal with long sequences, however sequences in image classification are typically not long enough to show the theoretical gains. On the contrary, it is more efficient to use shorter sequences instead. We believe that efficient mechanisms focusing on these shorter sequences are a fruitful area for future research.
@@ 5.6 Correlation of Metrics
We run a correlation analysis between the efficiency metrics in Table 2. The highest correlation of $0.89$ is between fine-tuning time and training memory. This suggests a common underlying factor or bottleneck, possibly related to the necessity of memory reads during training. Understanding this relationship can provide valuable insights into the factors influencing training efficiency. Intriguingly, the highest correlation coefficient involving a theoretical metric is $0.85$ between FLOPS and training memory, suggesting the possibility for a rough estimation. For other metrics, we reaffirm the limited reliability of estimating computational costs solely based on theoretical metrics [14, 4]. Consequently, assessing model efficiency in practice requires the empirical measurement of throughput and memory requirements.
@@ 6 Discussion & Conclusion
WhichTransformerto Favor?1. MainConstraint2. Phase3. Transformers to FavorSpeedMemoryInferenceTrainingInferenceTraining{ViT NextViT, ToMe, TokenLearner, }{ViT, TokenLearner, FocalNet}{ViT, NextViT, CoaT, CvT, }{ViT, ToMe, TokenLearner, NextViT, …}
Figure 6: Suggestions on Which Transformer to Favor in image classification. First, choose the main constraint (1.), speed, or memory, and the phase you want to optimize this constraint at (2.): Inference or training. Based on that, we provide recommendations on which transformer to favor (3.) from the taxonomy classes of Baseline, Sequence Reduction, Hybrid Attention, Non-Attention Shuffling and Sparse Attention. See the supplementary material for a full list of Pareto optimal models for each metric.
In this paper, we provide a thorough benchmark of efficiency oriented transformers for vision, facilitating a fair, competitive comparison of contributions from NLP and CV and revealing some unexpected trends. Our analysis shows ViT’s continued Pareto optimality across multiple metrics, reaffirming its effectiveness as a baseline, despite other models claiming to be more efficient. Results also prove that scaling the model size is, in most cases, more efficient than using higher resolution images. This goes against the trend of efficient models being evaluated using high resolution images [59, 37, 17, 61]. Comparing different efficiency metrics, we reaffirm the limitations of estimating computational costs solely based on theoretical metrics.
When addressing the question of which transformer to favor, our benchmark offers actionable insights in the form of models and strategies to use (see Figure 6). ViT remains the preferred choice overall. However, Token Sequence methods can become viable alternatives when speed and training efficiency are of importance. For scenarios with significant inference memory constraints, considering Hybrid CNN-attention models can prove advantageous.
Looking ahead, we posit that the establishment of this comparable, fair benchmark and analysis will drive substantial advancements in efficiency-oriented transformers for CV, benefitting researchers and practitioners making architectural decisions for resource-constrained environments.
Acknowledgements
This work was funded by the Carl-Zeiss Foundation under the Sustainable Embedded AI project (P2021-02-009).